OpenStudy (anonymous):

i got 11? Will give medal & fan!(:

OpenStudy (anonymous):

@azolotor

OpenStudy (anonymous):

I did 121 divided by 11

OpenStudy (stamp):

Equation of a circle is\[(x-h)^2+(y-k)^2=r^2\]where center is the point (h, k) and r is your radius.\[x^2+y^2=121\]means the circle has center (0, 0), and \[121=r^2\] Solve for r.

OpenStudy (stamp):

11 is the correct answer, but why did you "divide by 11"? The proper evaluation is to identify\[121=r^2\]and take the square root of each side. This leaves you with\[\sqrt 121=r\]\[r=11\]

OpenStudy (anonymous):

11

OpenStudy (anonymous):

I took the sq roo of it i knew that automaticaly

OpenStudy (stamp):

Ok, so ... good job? You are finished.